OpenStudy (anonymous):

Jake and Seth put some money into their money boxes every week. The amount of money (y), in dollars, in their money boxes after a certain amount of time (x), in weeks, is shown by the equations below: Jake: y = 5x + 85 Seth: y = 20x + 40 After how many weeks will Jake and Seth have the same amount of money in their money boxes and what will that amount be? 3 weeks, $100 3 weeks, $285 6 weeks, $150 100 weeks, $3

OpenStudy (anonymous):

@epoweritheta

OpenStudy (anonymous):

ok

OpenStudy (epoweritheta):

friend try to solve it on ur own , think

OpenStudy (epoweritheta):

@ajanijones dont just give the answer its against the CODE OF CONDUCT

OpenStudy (anonymous):

D: idk

OpenStudy (ajanijones):

So first go, 5 times 3 . 3 will represent x okay. So 5*3 = ? Then (?)+85 = ? Can you find out the ? for me.

OpenStudy (ajanijones):

Then for the second one you go, 20*x. x = 3 weeks So 20*3 = ? then (?) + 40 = what

OpenStudy (ajanijones):

so Whats 5*3 = ? and whats The answer you got for 5*3 + 85 = What?

OpenStudy (anonymous):

5 x 3 = 15 15 + 85 = 100 20 x 3 = 60 60 + 40 = 100

OpenStudy (ajanijones):

Good job :) So 3 weeks, $100 Is your answer.